What is the activity coefficient for each ion at the given ionic strength at 25 °C? Activity coefficients at various ionic strengths can be found in this table. SO 2 − 4 ( μ = 0.01 M) γ = Y 3 + ( μ = 0.1 M) γ = Pr 3 + ( μ = 0.001 M) γ = ( CH 3 CH 2 ) 3 NH + ( CH 3 CH 2 ) 3 NH + ( μ = 0.1 μ = 0.1 M) γ = γ =
Answer #1

1) SO 2 − 4   ( μ = 0.01 M), Activity coefficients   = 0.660

2) Y 3 + ( μ = 0.1 M) , Activity coefficients   = 0.18

3) Pr 3 + ( μ = 0.001 M) ,Activity coefficients   = 0.738

4) ( CH 3 CH 2 ) 3 NH + ( μ = 0.1 μ = 0.1 M) , Activity coefficients   = 0.79